# Non-equilibrium dynamic hyperuniform states

### v2

@@ I Introduction
In equilibrium disordered systems, to ensure a hyperuniform structure factor scaling $S(k\rightarrow 0)\sim k^{\alpha}$ with $\alpha>0$, as suggested in Ref. Lomba et al. (2018); Torquato (2018), the direct correlation function in Fourier space $\tilde{c}(k\rightarrow 0)\sim-\beta\tilde{v}(k\rightarrow 0)$, and the effective pair potential in Fourier space $\tilde{v}(k)$ should have a power-law scaling $\tilde{v}(k\rightarrow 0)\sim-\beta^{-1}k^{-\alpha}$, which is the repulsive interaction in “like-charged” particles, e.g., hyperuniform binary plasmas Lomba et al. (2017, 2018). Here $\beta=1/k_{B}T$ with $k_{B}$ and $T$ the Boltzmann constant and temperature of the system, respectively. Differently, in non-equilibrium systems, the ways to achieve hyperuniformity are not restricted by equilibrium constraints, and long-range interactions are not necessary. In recent decades, a growing number of non-equilibrium systems of short-range interacting particles have been found to form disordered hyperuniform states. Examples include jammed particles Zachary et al. (2011); Ricouvier et al. (2017); Dreyfus et al. (2015); Berthier et al. (2011), perfect glass Zhang et al. (2016), sheared particles Weijs et al. (2015); Weijs and Bartolo (2017); Tjhung and Berthier (2015); Wilken et al. (2020); Mitra et al. (2021), self-propelled particles Oppenheimer et al. (2022); Huang et al. (2021); Lei et al. (2019), interacting-diffusing particles Grassberger et al. (2016); Bertrand et al. (2019), sedimenting non-Brownian irregularly-shaped paricles Goldfriend et al. (2017), and non-reciprocally interacting robots Chen et al. (2024) etc. This research direction has attracted significant scientific attention lately. Here we review the recent progress of various approaches to obtain non-equilibrium dynamic disordered hyperuniform states theoretically, numerically and experimentally. The review is organized as follows: in Sec. II, we briefly introduce the random organization dynamics and relavent modified models, and review the hyperuniform states in the absorbing phase transition; in Sec. III, we review the hyperuniform fluids in various active particle systems, e.g., non-equilibrium hyperuniform fluids, chiral active fluids and circularly swimming algae systems, etc.; in Sec. IV, we show another type of approaches to achieve dynamical hyperuniform structures, i.e., the field patterns generated from phase separating systems via spinodal decomposition, like the Model-B, Model-H and active field theories; lastly, some concluding remarks on these dynamic hyperuniform states are given in Sec. V.
@@ II.1 Hyperuniformity in critical absorbing state of random organization model
Subsequently, it was found that structures at the critical absorbing transition in the random organization model exhibit abnormal fluctuations suppressed at large scales, indicating the presence of hyperuniformity Tjhung and Berthier (2015); Hexner and Levine (2015). As shown in Fig. 2(A), for the random organization model considered in Ref. Hexner and Levine (2015), as the strain amplitude $\gamma$ increases from the absorbing state towards the critical transition point $\gamma_{c}$, the hyperuniform scaling of density fluctuation $\sigma^{2}(l\rightarrow\infty)=\langle\rho(l)^{2}\rangle-\langle\rho(l)\rangle^{2}\sim l^{-2.45}$ gradually emerges at large length-scale $l$. Besides, in Ref. Hexner and Levine (2015), it was also confirmed that the absorbing transition belongs to the conserved directed percolation (CDP) universality class, which is a class of models sharing similar critical behaviors with universal critical exponents Rossi et al. (2000); Henkel et al. (2008). In order to confirm whether the models of the same universality class have the same hyperuniform structure characteristics at the critical point, the authors also investigated two other models within the CDP universality class, i.e., the Manna model and the conserved lattice gas (CLG) model. As shown in Fig. 2(B) and (C), for the CLG model in 2D, they calculated both the particle density fluctuation $\sigma^{2}(l)$ and the structure factor $S(k)$, and found the same hyperuniform fluctuation scaling $\sigma^{2}(l\rightarrow\infty)\sim l^{-2.45}$ as observed in the 2D RO model. Additionally, they also examined the structure factor near the critical point of the 2D conserved lattice gas model in a large system and found the scaling $S(k\rightarrow 0)\sim k^{0.45}$ at the critical absorbing transition point. Furthermore, as shown in Fig. 2(D) and (E), the authors analyzed 1D systems and found that the density fluctuation of the 1D RO and Manna models exhibited the same hyperuniform scaling $\sigma^{2}(l\rightarrow\infty)\sim l^{-1.425}$ at the critical point. This suggests that the hyperuniform scaling observed at the critical point of the CDP universality class depends on the dimensionality of the system, and the critical hyperuniform scaling at the absorbing transition was found to be $S(k\rightarrow 0)\sim k^{0.25}$ in 3D CLG model Hexner and Levine (2015). Later, it was confirmed that the critical hyperuniformity observed at the absorbing transition is robust against the change of particle shape in the RO model Ma and Torquato (2019). Below the critical point of absorbing transition $\rho_{c}$, the system eventually evolves into an absorbing state, which is hyperuniform up to a length-scale that diverges at $\rho_{c}$. Hexner et al. found that by adding extra weak noises to structures in the absorbing state, one can enhance the hyperuniformity of the absorbing state Hexner et al. (2017).
@@
Moreover, in a subsequent study Wilken et al. (2021), the authors found that this modified RO model can effectively establish a dynamical approach towards the random close packing (RCP). Comparing to the original random organization model proposed in Ref. Corté et al. (2008), they investigated the zero-strain limit ($\gamma=0$) and found that the biased random organization dynamics resulted in a higher critical packing fraction. Remarkably, they found that this critical packing fraction is very close to the volume fraction of random close packing $\phi_{\rm RCP}$ Ni et al. (2013). Furthermore, they demonstrated that the configurations obtained from the biased random organization critical point appeared to be structurally identical to RCP configurations obtained by other protocols. These similarities included isostatic coordination ($Z=6$), a cusp in the pair correlation function $g(r)$, and the structure factor scaling of $S(k\rightarrow 0)\sim k^{0.25}$ in 3D. Afterwards, it was found that the RCP structures obtained by this dynamic approach are hyperuniform only in $d<4$ dimensions Wilken et al. (2023a).
@@ II.2 Hyperuniformity in self-organized criticality based on random organization model
In the previous section, we have mentioned recent studies on the hyperuniformity at the critical point of absorbing phase transitions. In Ref. Wang et al. (2018), Wang et al. numerically investigated an intriguing non-fine-tuned hyperuniform system based on self-organized critical (SOC) states in sheared non-Brownian suspension experiments proposed by Corté et al. in Ref. Corté et al. (2009). Specifically, based on the random organization dynamics in Fig. 1(A), the SOC random organization model introduced an additional sedimentation velocity $v_{s}$ acting on all particles towards a specific direction to mimic the gravity-induced sedimentation in cyclically sheared non-Brownian suspensions. The authors defined the horizontal $x$-axis as the shear flow direction and the vertical $y$-axis as the sedimentation direction, as shown in Fig. 4(A). In computer simulations, $N$ particles of diameter $d$ were placed in a square box of width $L$, with an area fraction $\phi=N\pi(d/2)^{2}/L^{2}$. Periodic boundary condition is applied in $x$-axis, and a hard impenetrable wall is placed at the bottom of the box ($y=0$). Additionally, a linear density of particles was defined as $\kappa\equiv N/L$ along the $x$-axis. In this system, two competing dynamics are at play: sedimentation dynamics and shear-induced random organization dynamics. Sedimentation dynamics tend to concentrate particles towards the bottom of the box in $y$ direction, while the repulsive pair-interactions driven by random organization dynamics tend to push particles away from each other against sedimenting to the bottom. As demonstrated in Ref. Corté et al. (2009); Wang et al. (2018), due to these competing dynamics, the system automatically self-organizes into a uniform or non-uniform steady state, as shown in Fig. 4(B), depending on the sedimentation speed $v_{s}$. At low sedimentation speed $v_{s}$, particles are more likely to self-organize into a uniform distribution in the $y$ direction with a steady-state height $h^{\infty}$, while this is not observed at high sedimentation speed where the sedimentation effect dominates so that more particles tend to stay near the bottom. Remarkably, in the low $v_{s}$ case, it was found that the constant mean area fraction, defined as $\bar{\phi}^{\infty}\equiv\pi d^{2}\kappa/4\bar{h}^{\infty}$, converges to the critical volume fraction of the random organization model at the same strain amplitude $\gamma_{0}$.
@@
In the absence of thermal diffusion, the area fraction at the bottom wall would increase due to sedimentation from $\phi_{c}$ to $2\phi_{c}$ in a time $\tau_{s}=\xi/v_{s}$, where $\xi=\sqrt{\pi d^{2}/4\phi_{c}}$ is the mean interparticle spacing. Meanwhile, shearing would also induce diffusion and redistribute the increased $\phi$ due to sedimentation to a uniform steady state with a suspension height $h^{\infty}$, characterized by a typical time scale $\tau={h^{\infty}}^{2}/4D$, where $D$ is the diffusivity under the strain amplitude $\gamma_{0}$ and the area fraction $2\phi_{c}$. Thus, it was found that the competing dynamics can be well captured by a dimensionless factor $A=\tau_{D}/\tau_{s}=\left(\pi/\phi_{c}\right)^{3/2}d^{3}\kappa^{2}v_{s}/(4D)$ Corté et al. (2009). In the limit $A\ll 1$, the authors found that the system evolves into a uniform steady state with a self-organized volume fraction $\phi_{c}$. In Ref. Wang et al. (2018), for better data collapsing, by taking $\tau^{\prime}_{D}$ and $\tau^{\prime}_{s}$ as the timescales for particle transport over the same length scale $h_{c}$ the critical height of the bed of particles, Wang et al. defined a new dimensionless factor
@@ II.3 Hyperuniformity in active states of random organization model
In Ref. Hexner and Levine (2017), Hexner and Levine studied variants of models based on the random organization and the Manna model. Comparing to the original model, they introduced the center-of-mass conservation (CMC) into the system and found that this modification led to new dynamic hyperuniform states. As shown in the left panel of Fig. 5(A), pairs of active particles are given displacements along the axis connecting their centers, with the same amplitudes chosen from a uniform distribution in the range $[0,\sigma]$, where $\sigma$ is the particle diameter, ensuring the center-of-mass conservation. In contrast to the original random organization model Corté et al. (2008); Hexner and Levine (2015), the CMC version of the model exhibits less randomness in active particle diffusion because the direction of each collision is deterministic rather than randomly chosen. In the study of the 2D zero-strain case, the structure factors of systems at several particle densities $\rho$ are shown in Fig. 5(B). They observed that the structure factor exhibited the same scaling $S(k\rightarrow 0)\sim k^{0.45}$ as the original RO model at the critical point of the absorbing-state phase transition. Moreover, they found that the active states exhibited a new dimension-irrelevant structure factor scaling $S(k\rightarrow 0)\sim k^{2}$, which indicates that the center-of-mass conservation suppresses long-range fluctuations in the active state and induces strong long-ranged hyperuniform correlations. Additionally, they studied a CMC version of the Manna model, where particles in active sites could redistribute in opposite directions to adjacent neighboring sites, and the structure factor shows the same hyperuniform scaling in the corresponding critical state and active states, as shown in Fig. 5(C).
@@
where $D=\omega a^{2}$ is the intrinsic diffusion constant and $a$ is the lattice constant, the Gaussian white noise $\left\langle\eta(x,t)\eta\left(x^{\prime},t^{\prime}\right)\right\rangle=\delta\left(x-x^{\prime}\right)\delta\left(t-t^{\prime}\right)$ with the noise amplitude $A=\sqrt{\omega_{0}}a^{2}$. To understand this quadratic noise term $\partial_{xx}\eta$, the authors considered the system in the continuum limit, and the particle number conservation could be written as
@@
To ensure $\partial R_{\alpha}=0$, $J_{\alpha}$ in the last term of Eq. 11 should be written in the divergence form as $J_{\alpha}=-\nabla\cdot\sigma_{\alpha}$, which leads Eq. 10 to $\delta_{t}n=\nabla\cdot\nabla\cdot\sigma_{\alpha}$. In 1D, $\sigma$ could be written as $\sigma=f(n)+g(n)\eta$, where $f(n)$ is deterministic and $g(n)\eta$ with Gaussian white noise $\eta$ represents the stochastic motion. Thus, in Eq. 10, the term accounting for stochastic motion should be written in the quadratic divergence form $\partial_{xx}\eta$. This center-of-mass conserved active state leads to a new dynamic hyperuniform state and further inspires a number of non-equilibrium hyperuniform fluid systems, e.g., chiral active fluids and non-equilibrium driven-dissipative hyperuniform fluids Lei et al. (2019); Lei and Ni (2019); Tjhung and Berthier (2017), which are discussed in the next section.

### v3

@@ I Introduction
In equilibrium disordered systems, to ensure a hyperuniform structure factor scaling $S(k\rightarrow 0)\sim k^{\alpha}$ with $\alpha>0$, as suggested in Ref. Lomba et al. (2018); Torquato (2018), the direct correlation function in Fourier space $\tilde{c}(k\rightarrow 0)\sim-\beta\tilde{v}(k\rightarrow 0)$, and the effective pair potential in Fourier space $\tilde{v}(k)$ should have a power-law scaling $\tilde{v}(k\rightarrow 0)\sim-\beta^{-1}k^{-\alpha}$, which is the repulsive interaction in “like-charged” particles, e.g., hyperuniform binary plasmas Lomba et al. (2017, 2018). Here $\beta=1/k_{B}T$ with $k_{B}$ and $T$ the Boltzmann constant and temperature of the system, respectively. Differently, in non-equilibrium systems, the ways to achieve hyperuniformity are not restricted by equilibrium constraints, and long-range interactions are not necessary. In recent decades, a growing number of non-equilibrium systems of short-range interacting particles have been found to form disordered hyperuniform states. Examples include jammed particles Zachary et al. (2011); Ricouvier et al. (2017); Dreyfus et al. (2015); Berthier et al. (2011), perfect glass Zhang et al. (2016), sheared particles Weijs et al. (2015); Weijs and Bartolo (2017); Tjhung and Berthier (2015); Wilken et al. (2020); Mitra et al. (2021), self-propelled particles Oppenheimer et al. (2022); Huang et al. (2021); Lei et al. (2019), interacting-diffusing particles Grassberger et al. (2016); Bertrand et al. (2019), sedimenting non-Brownian irregularly-shaped paricles Goldfriend et al. (2017), non-reciprocally interacting robots Chen et al. (2024), driven-diffusive particles with exclusion process Jack et al. (2015); Karevski and Schütz (2017); Garrahan and Pollmann (2024) and quantum systems Torquato et al. (2008); Carollo et al. (2017); Sakai et al. (2022) etc. This research direction has attracted significant scientific attention lately. Here we review the recent progress of various approaches to obtain non-equilibrium dynamic hyperuniform states in classical many-particle systems theoretically, numerically and experimentally. The review is organized as follows: in Sec. II, we briefly introduce the random organization dynamics and relavent modified models, and review the hyperuniform states in the absorbing phase transition; in Sec. III, we review the hyperuniform fluids in various active particle systems, e.g., non-equilibrium hyperuniform fluids, chiral active fluids and circularly swimming algae systems, etc.; in Sec. IV, we show another type of approaches to achieve dynamical hyperuniform structures, i.e., the field patterns generated from phase separating systems via spinodal decomposition, like the Model-B, Model-H and active field theories; lastly, some concluding remarks on these dynamic hyperuniform states are given in Sec. V.
@@ II.1 Hyperuniformity in critical absorbing state of random organization model
Subsequently, it was found that structures at the critical absorbing transition in the random organization model exhibit abnormal fluctuations suppressed at large scales, indicating the presence of hyperuniformity Tjhung and Berthier (2015); Hexner and Levine (2015). As shown in Fig. 2(A), for the random organization model considered in Ref. Hexner and Levine (2015), as the strain amplitude $\gamma$ increases from the absorbing state towards the critical transition point $\gamma_{c}$, the hyperuniform scaling of density fluctuation $\sigma^{2}(l\rightarrow\infty)=\langle\rho(l)^{2}\rangle-\langle\rho(l)\rangle^{2}\sim l^{-2.45}$ gradually emerges at large length-scale $l$. Besides, in Ref. Menon and Ramaswamy (2009), it was confirmed that the absorbing state transition for random organization model belongs to the conserved directed percolation (CDP) universality class, which is a class of models sharing similar critical behaviors with universal critical exponents Rossi et al. (2000); Henkel et al. (2008). In order to confirm whether other models in the same universality class exhibit the same hyperuniform structural characteristics at the critical point, Ref. Hexner and Levine (2015) also investigated two other models within the CDP universality class, namely, the Manna model and the conserved lattice gas (CLG) model. As shown in Fig. 2(B) and (C), for the CLG model in 2D, they calculated both the particle density fluctuation $\sigma^{2}(l)$ and the structure factor $S(k)$, and found the same hyperuniform fluctuation scaling $\sigma^{2}(l\rightarrow\infty)\sim l^{-2.45}$ as observed in the 2D RO model. Additionally, they also examined the structure factor near the critical point of the 2D conserved lattice gas model in a large system and found the scaling $S(k\rightarrow 0)\sim k^{0.45}$ at the critical absorbing transition point. Furthermore, as shown in Fig. 2(D) and (E), the authors analyzed 1D systems and found that the density fluctuation of the 1D RO and Manna models exhibited the same hyperuniform scaling $\sigma^{2}(l\rightarrow\infty)\sim l^{-1.425}$ at the critical point. This suggests that the hyperuniform scaling observed at the critical point of the CDP universality class depends on the dimensionality of the system, and the critical hyperuniform scaling at the absorbing transition was found to be $S(k\rightarrow 0)\sim k^{0.25}$ in 3D CLG model Hexner and Levine (2015). Later, it was confirmed that the critical hyperuniformity observed at the absorbing transition is robust against the change of particle shape in the RO model Ma and Torquato (2019). Below the critical point of absorbing transition $\rho_{c}$, the system eventually evolves into an absorbing state, which is hyperuniform up to a length-scale that diverges at $\rho_{c}$. Besides, Hexner et al. found that by adding extra weak noises to structures in the absorbing state, for example, reactivating the system by giving small thermal-noise-like displacements to all particles and waiting for the noise-induced activity to vanish, one can enhance the hyperuniformity of the absorbing state Hexner et al. (2017). Moreover, Mari et al. modified the original RO dynamics with mediated interactions where passive particles undergo total-activity related diffusions to mimic the hydrodynamic effect in realistic systems, and observed that the hyperuniformity at critical absorbing state is destroyed at large length-scales Mari et al. (2022).
@@
Moreover, in a subsequent study Wilken et al. (2021), it was found that this modified RO model can effectively establish a dynamical approach towards the random close packing (RCP). Comparing to the original random organization model proposed in Ref. Corté et al. (2008), they investigated the zero-strain limit ($\gamma=0$) and found that the biased random organization dynamics resulted in a higher critical packing fraction. Remarkably, they found that this critical packing fraction is very close to the volume fraction of random close packing $\phi_{\rm RCP}$ Ni et al. (2013). Furthermore, they demonstrated that the configurations obtained from the biased random organization critical point appeared to be structurally identical to RCP configurations obtained by other protocols. These similarities included isostatic coordination ($Z=6$), a cusp in the pair correlation function $g(r)$, and the structure factor scaling of $S(k\rightarrow 0)\sim k^{0.25}$ in 3D. Afterwards, it was found that the RCP structures obtained by this dynamic approach are hyperuniform only in $d<4$ dimensions Wilken et al. (2023a).
@@ II.2 Hyperuniformity in self-organized criticality based on random organization model
In the previous section, we have mentioned recent studies on the hyperuniformity at the critical point of absorbing phase transitions. In Ref. Wang et al. (2018), Wang et al. numerically investigated an intriguing non-fine-tuned hyperuniform system based on self-organized critical (SOC) states in sheared non-Brownian suspension experiments proposed by Corté et al. in Ref. Corté et al. (2009). Specifically, based on the random organization dynamics in Fig. 1(A), the SOC random organization model introduced an additional sedimentation velocity $v_{s}$ acting on all particles towards a specific direction to mimic the gravity-induced sedimentation in cyclically sheared non-Brownian suspensions. The authors defined the horizontal $x$-axis as the shear flow direction and the vertical $y$-axis as the sedimentation direction, as shown in Fig. 4(A). In computer simulations, $N$ particles of diameter $d$ were placed in a square box of width $L$, with an area fraction $\phi=N\pi(d/2)^{2}/L^{2}$. Periodic boundary condition is applied in $x$-axis, and a hard impenetrable wall is placed at the bottom of the box ($y=0$). Additionally, a linear density of particles was defined as $\kappa\equiv N/L$ along the $x$-axis. In this system, two competing dynamics are at play: sedimentation dynamics and shear-induced random organization dynamics. Sedimentation dynamics tend to concentrate particles towards the bottom of the box in $y$ direction, while the effective repulsive pair-interactions via random organization dynamics tend to push particles away from each other against sedimenting to the bottom. As demonstrated in Ref. Corté et al. (2009); Wang et al. (2018), due to these competing dynamics, the system automatically self-organizes into a uniform or non-uniform steady state, as shown in Fig. 4(B), depending on the sedimentation speed $v_{s}$. At low sedimentation speed $v_{s}$, particles are more likely to self-organize into a uniform distribution in the $y$ direction with a steady-state height $h^{\infty}$, while this is not observed at high sedimentation speed where the sedimentation effect dominates so that more particles tend to stay near the bottom. Remarkably, in the low $v_{s}$ case, it was found that the constant mean area fraction, defined as $\bar{\phi}^{\infty}\equiv\pi d^{2}\kappa/4\bar{h}^{\infty}$, converges to the critical volume fraction $\phi_{c}$ of the random organization model at the same strain amplitude $\gamma_{0}$.
@@
In the absence of thermal diffusion, the area fraction at the bottom wall would increase due to sedimentation from $\phi_{c}$ to $2\phi_{c}$ in a time $\tau_{s}=\xi/v_{s}$, where $\xi=\sqrt{\pi d^{2}/4\phi_{c}}$ is the mean interparticle spacing. Meanwhile, shearing would also induce diffusion and redistribute the increased $\phi$ due to sedimentation to a uniform steady state with a suspension height $h^{\infty}$, characterized by a typical time scale $\tau_{D}={h^{\infty}}^{2}/4D$, where $D$ is the diffusivity under the strain amplitude $\gamma_{0}$ and the area fraction $2\phi_{c}$. Thus, it was found that the competing dynamics can be well captured by a dimensionless factor $A=\tau_{D}/\tau_{s}=\left(\pi/\phi_{c}\right)^{3/2}d^{3}\kappa^{2}v_{s}/(4D)$ Corté et al. (2009). In the limit $A\ll 1$, the authors found that the system evolves into a uniform steady state with a self-organized volume fraction $\phi_{\infty}=\phi_{c}$, and $\phi_{\infty}/\phi_{c}$ can be collapsed as a function of $A$ for various linear densities $\kappa$ and strain amplitudes $\gamma_{0}$ at small $A$. Later, in Ref. Wang et al. (2018), Wang et al. found the dimensionless factor $A$ can not be used to describe high sedimentation speed case ($A\gg 1$), as it results in poor data collapse for different linear densities $\kappa$ with increasing $A$. For better data collapsing, the authors proposed that the timescales for diffusion and sedimentation should instead be considered over the same length-scale, and defined a new dimensionless factor $\bar{A}$. By taking $\tau^{\prime}_{D}$ and $\tau^{\prime}_{s}$ as the new timescales for particle transport over the critical height of the bed of particles, $h_{c}$, Wang et al. defined a new dimensionless factor
@@ II.3 Hyperuniformity in active states of random organization model
In Ref. Hexner and Levine (2017), Hexner and Levine studied variants of models based on the random organization and the Manna model. Comparing to the original model, they introduced the center-of-mass conservation (CMC) into the system and found that this modification led to new dynamic hyperuniform states. As shown in the left panel of Fig. 5(A), pairs of active particles are given displacements along the axis connecting their centers, with the same amplitudes chosen from a uniform distribution in the range $[0,\sigma]$, where $\sigma$ is the particle diameter, ensuring the center-of-mass conservation. Comparing to the other modified random organization model we have discussed in Section II.A Wilken et al. (2020), the CMC version of the RO model can be regarded as setting $\epsilon_{r}=0$ only retaining the reciprocal repulsions. This adjustment enforces the system to exhibit less randomness in active particle diffusion as the direction of each collision is deterministic rather than randomly chosen. In the study of the 2D zero-strain case, the structure factors of systems at several particle densities $\rho$ are shown in Fig. 5(B). They observed that the structure factor exhibited the same scaling $S(k\rightarrow 0)\sim k^{0.45}$ as the original RO model at the critical point of the absorbing-state phase transition. Moreover, they found that the active states exhibited a new dimension-irrelevant structure factor scaling $S(k\rightarrow 0)\sim k^{2}$, which indicates that the center-of-mass conservation suppresses long-range fluctuations in the active state and induces strong long-ranged hyperuniform correlations. Additionally, they studied a CMC version of the Manna model, where particles in active sites could redistribute in opposite directions to adjacent neighboring sites, and the structure factor shows the same hyperuniform scaling in the corresponding critical state and active states, as shown in Fig. 5(C).
@@
where $D=\omega a^{2}$ is the intrinsic diffusion constant and $a$ is the lattice constant, the Gaussian white noise $\left\langle\eta(x,t)\eta\left(x^{\prime},t^{\prime}\right)\right\rangle=\delta\left(x-x^{\prime}\right)\delta\left(t-t^{\prime}\right)$ with the noise amplitude $A=\sqrt{\omega}a^{2}$. To understand this quadratic noise term $\partial_{xx}\eta$, the authors considered the system in the continuum limit, and the particle number conservation could be written as
@@
To ensure $\partial_{t}R_{\alpha}=0$, $J_{\alpha}$ in the last term of Eq. 11 should be written in the divergence form as $J_{\alpha}=-\nabla\cdot\sigma_{\alpha}$, which leads Eq. 10 to $\delta_{t}n=\nabla\cdot\nabla\cdot\sigma_{\alpha}$. In 1D, $\sigma$ could be written as $\sigma=f(n)+g(n)\eta$, where $f(n)$ is deterministic and $g(n)\eta$ with Gaussian white noise $\eta$ represents the stochastic motion. Thus, in Eq. 10, the term accounting for stochastic motion should be written in the quadratic divergence form $\partial_{xx}\eta$. This center-of-mass conserved active state leads to a new dynamic hyperuniform state and further inspires a number of non-equilibrium hyperuniform fluid systems, e.g., chiral active fluids and non-equilibrium driven-dissipative hyperuniform fluids Lei et al. (2019); Lei and Ni (2019); Tjhung and Berthier (2017), which are discussed in the next section.